OpenStudy (anonymous):

Find the solution to the equation. e^5x=25

OpenStudy (freckles):

e^(5x)=25 take ln( ) of both sides

OpenStudy (freckles):

This is so you can bring down the (5x) part

OpenStudy (anonymous):

okay, i have no idea how to to this equation at all! Could you help?

OpenStudy (freckles):

Well have you taken the ln( ) of both sides as I suggested ?

OpenStudy (freckles):

\[e^{5x}=25 \\ \ln(e^{5x})=\ln(25)\] now just use power rule on the left hand side bring down the 5x in front of the ln(e) and recall ln(e)=1

OpenStudy (freckles):

can you tried to do that? bring down the 5x part in front of the ln(e)

OpenStudy (anonymous):

you had a math processing error up there^

OpenStudy (freckles):

try refreshing i don't see one on my end

OpenStudy (freckles):

But if you can't read it: 1st step: take ln( ) of both sides 2nd step: use power rule so you bring the 5x thing 3rd step: use that ln(e) is 1 4th step: just divide both sides by what x is being multiplied by.

OpenStudy (freckles):

if you want to type what you have so far I can check it